OpenStudy (anonymous):

Let f(x,y) = (−(4x+y))5. Then [(∂2f)/(∂x∂y)] = ??? [(∂3f)/(∂x∂y∂x)] = ??? [(∂3f)/(∂x2∂y)] = ???

myininaya (myininaya):

\[f=(-(4x+y))^5?\]

OpenStudy (anonymous):

yes

myininaya (myininaya):

\[f=-(4x+y)^5 \] \[f_x=-5(4x+y)^{5-1}(4)=-20(4x+y)^4\]

myininaya (myininaya):

\[f_{xy}=-20 \cdot 4(4x+y)^{4-1}(1)=-80(4x+y)^3\]

myininaya (myininaya):

all you have to do is treat everything else like a constant except the one you are differentiating with respect to

myininaya (myininaya):

\[f_{xyx}=-80(3)(4x+y)^{3-1}(4)\]

myininaya (myininaya):

you should try the last one

myininaya (myininaya):

i will be happy to check it

OpenStudy (anonymous):

why multiply that last one by 4 ?!

myininaya (myininaya):

chain rule

OpenStudy (anonymous):

so -900(4x+y)^2 ?

myininaya (myininaya):

partial derivative with respect to x of 4x+y is 4 parital derivative with respect to y of 4x+y is 1

OpenStudy (anonymous):

i know that ! but my question is how is that thing works !? [(∂^3f)/(∂x∂y∂x)]

myininaya (myininaya):

?

OpenStudy (anonymous):

i know its higher order but the signs for it are hella confusing

OpenStudy (anonymous):

so in the last one i do it considering x twice then with y one !?

myininaya (myininaya):

yep

myininaya (myininaya):

i already did f_x now you have to find f_xx then f_xxy

OpenStudy (anonymous):

ok cool thanks

myininaya (myininaya):

i like this notation more than the one above it looks less messy i think lol